Please help me with my probability assignment.Could you please give me a complete solution to these questions.
1. A jar has 4 red marbles, 5 green marbles and 7 blue marbles. A marble is drawn, observed and then replaced. Another marble is then drawn. Calculate:
a.) The probability that both were red marbles
b.) The probability that there was one red and one green marble
c.) The probability that neither was a blue marble
2. In a group of 100 students, 55 study Mathematical Methods and 40 study Physics. If 12 study neither subject, how many study both?

Answer by stanbon(75887)

1. A jar has 4 red marbles, 5 green marbles and 7 blue marbles. A marble is drawn, observed and then replaced. Another marble is then drawn. Calculate:
a.) The probability that both were red marbles
The draws are independent because you replaced the 1st draw.
P(red and red) = P(red)*P(red) = (4/16)(4/16) = 1/16
--------------------------------------------------------
b.) The probability that there was one red and one green marble
P(red and green) = (4/16)(5/16)
-------------------------------------------
c.) The probability that neither was a blue marble
P(not blue and not blue) = (9/16)^2
------------------------------------------
2. In a group of 100 students, 55 study Mathematical Methods and 40 study Physics. If 12 study neither subject, how many study both?
---
Draw a Venn Diagram.
If 12 study neither, 88 study MM or Phy or both subjects.
With the information given, you cannot tell how many studied both.
======================================================
Cheers,
Stan H.